OpenStudy (anonymous):

Pls help WILL FAN AND MEDAL Please Help!! Evaluate. cube root of 5 multiplied by square root of 5 over cube root of 5 to the power of 5

OpenStudy (anonymous):

5 to the power of negative 1 over 6 5 to the power of 3 over 2 5 to the power of 5 over 2 5 to the power of negative 5 over 6

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@Sushi121212

OpenStudy (anonymous):

i think its the last one

OpenStudy (anonymous):

but i need help to undestand it

OpenStudy (anonymous):

@bibby

OpenStudy (tkhunny):

Have you considered writing all the elements as an exponent on the common base, 5? \(\sqrt{5} = 5^{1/2}\) \(\sqrt[3]{5} = 5^{1/3}\) \(\sqrt{5}\cdot\sqrt[3]{5} = 5^{1/2}\cdot 5^{1/3} = 5^{1/2 + 1/3} = 5^{3/6 + 2/6} = 5^{5/6} = \left(\sqrt[6]{5}\right)^{5}\) Let's see what you get out of the denominator.